OpenStudy (anonymous):

Find the slope-intercept form of the equation of a line with slope 1/2 containing point (9,9). Question 8 options: x=1/2y+4 1/2 y=1/2x+13 1/2 y=4 1/2x+1/2 y=1/2x+4 1/2

OpenStudy (accessdenied):

put it into point-slope form and solve for y. y - y1 = m(x - x1) for a given point (x1, y1) and slope m

hero (hero):

y - 9 = .5(x - 9) y - 9 = .5x - 4.5 y = .5x - 4.5 + 9 y = .5x + 4.5 y = (x+9)/2

hero (hero):

last option by the way